1 The Croatian language

The file croatian.dtx^{1} defines all the language definition macros for the Croatian language.

For this language currently no special definitions are needed or available.

The macro \LdfInit takes care of preventing that this file is loaded more than once, checking the category code of the @ sign, etc.

1 ⟨∗code⟩

 2 \LdfInit{croatian}\captionscroatian

When this file is read as an option, i.e. by the \usepackage command, croatian will be an ‘unknown’ language in which case we have to make it known. So we check for the existence of \l@croatian to see whether we have to do something here.

 3 \ifx\l@croatian\@undefined
 4  \@nopatterns{Croatian}
 5  \adddialect\l@croatian0\fi

The next step consists of defining commands to switch to (and from) the Croatian language.

\captionscroatian The macro \captionscroatian defines all strings used in the four standard documentclasses provided with L^{A}TEX.

 6 \addto\captionscroatian{%
 7 \def\prefacename{Predgovor}%
 8 \def\refname{Literatura}%
 9 \def\abstractname{Sa\v{z}etak}%
10 \def\bibname{Bibliografija}%
11 \def\chaptername{Glava}%
12 \def\appendixname{Dodatak}%
13 \def\contentsname{Sadr\v{z}aj}%
14 \def\listfigurename{Slike}%
15 \def\listtablename{Tablice}%
16 \def\indexname{Indeks}%
17 \def\figurename{Slika}%
18 \def\tablename{Tablica}%
19 \def\partname{Dio}%
20 \def\enclname{Prilozi}%
21 \def\ccname{Kopije}%
22 \def\headtoname{Prima}%
23 \def\pagename{Strana}%
24 \def\seename{Vidi}%
25 \def\alsoname{Vidi tako\dj er}%
26 \def\proofname{Dokaz}%
27 }%

\datecroatian The macro \datecroatian redefines the command \today to produce Croatian dates.

28 \def\datecroatian{%
29 \def\today{\number\day .~\ifcase\month\or
30  sijev{c}nja\or velja\v{c}e\or o\v{z}ujka\or travnja\or svibnja\or

1The file described in this section has version number v1.3g and was last revised on 1996/12/23. A contribution was made by Alan Paić (paica@cernvm.cern.ch).

31  lipnja\or srpnja\or kolovoza\or rujna\or listopada\od studenog\or
32  prosinca\fi \space \number\year}}

\extrascroatian \noextrascroatian The macro \extrascroatian will perform all the extra definitions needed for the Croatian language. The macro \noextrascroatian is used to cancel the actions of \extrascroatian. For the moment these macros are empty but they are defined for compatibility with the other language definition files.

33 \addto\extrascroatian{}
34 \addto\noextrascroatian{}

The macro \ldf@finish takes care of looking for a configuration file, setting the main language to be switched on at \begin{document} and resetting the category code of @ to its original value.

35 \ldf@finish{croatian}

36 ⟨/code⟩